solve the equation d²y/dt² + 2dy/dt -3y = sin t , given that y = dy/dt = 0 when t = 0
1
Expert's answer
2020-12-29T15:19:38-0500

"\frac{d^2y}{dt^2}+2\frac{dy}{dt}-3y=\sin t"

The corresponding homogeneous equation "y''+2y'-3y=0" has the characteristics equation "m^2+2m-3=0"

"(m-1)(m+3)=0\\
m=1\text{ or } m=-3"

"Y_c=C_1e^{1.t}+C_2e^{-3t}\\
Y_c=C_1e^t+C_2e^{-3t}"

The non-homogeneous term is "g(t)=\sin t" . We choose "Y_p=A\cos t+ B\sin t"

"Y'_p=B\cos t-A\sin t, Y''_p=-(A\cos t+B\sin t)" . Put these into the equation. We have;

"-(A\cos t+B\sin t)+2(B\cos t-A\sin t)-3(A\cos t+B\sin t)=\sin t\\
 (-4A+2B)\cos t+(-2A-4B)\sin t=\sin t"

Compare both sides

"-4A+2B=0.................(1)\\
-2A-4B=1................(2)"

From (1), "A=\frac{B}{2}." Put this into (2)

"-5B=1,B=-\frac{1}{5}\\
A=-\frac{1}{10}"

Therefore,"Y_p=-\frac{1}{10}\cos t-\frac{1}{5}\sin t"

The general solution "y=Y_c+Y_p"


"y=C_1e^t+C_2e^{-3t}-\frac{1}{10}\cos t-\frac{1}{5}\sin t"


But, "y=y'=0" when "t=0" . This implies that,

"0=C_1e^0+C_2e^{-3.0}-\frac{1}{10}\cos 0-\frac{1}{5}\sin 0\\
0=C_1+C_2-\frac{1}{10}\\
\frac{1}{10}=C_1+C_2....................(3)"


"y'=C_1e^t-2C_3e^{-3t}+\frac{1}{10}\sin t-\frac{1}{5}\cos t\\
0=C_1e^0-3C_2e^{-3.0}+\frac{1}{10}\sin 0-\frac{1}{5}\sin 0\\
0=C_1-3C_2-\frac{1}{5}\\
C_1-3C_2=\frac{1}{5}....................(4)"

Subtract (4) from (3)

"4C_2=\frac{1}{10}-\frac{1}{5}=-\frac{1}{10},C_2=-\frac{1}{40}"

Put this into (3)

"C_1=\frac{1}{10}+\frac{1}{40}=\frac{1}{8}"


Hence, "y=\frac{1}{8}e^t-\frac{1}{40}e^{-3t}-\frac{1}{10}\cos t-\frac{1}{5}\sin t"